# Automatic Hardware Pragma Insertion in High-Level Synthesis: A Non-Linear Programming Approach

### v2

@@ 4.3. Summary
By composing all the theorems, this allows us to end up with the final latency lower bound of the program which is presented in theorems B.20 for the computation and B.21 for the computation and communication.
@@
In Listing 5, let’s denote the loop body with loop $L0$ and statement $S0$ as $LB0$, and the one with loops $L1$ and $L2$, along with statements $S1$, $S2$, and $S3$, as $LB1$.
For $LB0$, where the loop is unrolled with a factor $uf$ and no dependency exists between loop iterations, the latency is bounded by $\frac{N}{uf}\times 1$, assuming the initialization takes at least one cycle.
Regarding $LB1$, loop $L2$ is fully unrolled, with no dependencies between iterations for statement $S2$. As there is no dependency between $S2$ and $S3$, they can be executed in parallel, and the latency of the unrolled part equals the critical path, which is the maximum between the unrolled parts of $S2$ and $S3$ in this scenario. However, for statement $S3$, despite the absence of dependencies among multiplication operations enabling parallel execution, a tree reduction is involved. Given the pipelined nature of loop $L1$, functioning as a reduction for statement $S2$, we observe $II>IL_{+}$, where $IL_{+}$ denotes the latency of the reduction operation. Consequently, the latency of $LB1$ is bounded by $(\max(IL_{+}+IL_{*},IL_{+}\log_{2}(N)\times IL_{+})+IL_{+}\times(M-1))$.
@@
Given available resource $DSP_{avail}$ and a program which contains a loop body $L$ with a computation latency $Lat_{L}^{computation}$ and a communication latency $Lat_{L}^{communication}$.
@@
$Lat_{L}=Lat_{L}^{computation}+Lat_{L}^{communication}$
5. Non-Linear Formulation for Pragma Insertion
Let $\mathcal{P}$ be the set of different possible pipeline configurations. Let $\forall p\in\mathcal{P}$ define $\mathcal{L}^{p}_{pip}$ the set of loops pipelined and $\forall l\in\mathcal{L}^{p}_{pip}$, $\mathcal{L}^{p}_{under\_pip_{l}}$ the set of loops under a loop pipelined and $\mathcal{L}^{p}_{above\_pip_{l}}$ the set of loops above the loop pipelined $l$. Let $\forall s\in\mathcal{S}$ define the set of nested loops which iterate the statement $s$, $\mathcal{L}_{s}$. $\forall a\in\mathcal{A}$ and for $d$ a dimension of the array $a$, let $\mathcal{C}_{a_{d}}$ be the set of loops which iterates the array $a$ at the dimension $d$. $\forall l\in\mathcal{L}$, let designate $d_{l}$ the maximum dependency distance of the loop $l$. And let $II_{s}$ be the II of the loop pipelined for the statement $s$.
The II for each loop, the dependencies, the properties of the loops, the TC, the iteration latency of the parallel operations and the reduction operations and the number of DSPs per operation per statements are computed at compile time with PolyOpt-HLS (Pouchet et al., 2013b) and used as constants in the NLP problem.
5.1. Variables
Set Description
$\mathcal{L}$ the set of loops
$\mathcal{A}$ the set of arrays
$\mathcal{S}$ the set of statements
$\mathcal{O}_{s}$ the list of operations of the statements $s$
$\mathcal{O}_{s_{par}}$ the operation which can be done in parallel, i.e., does not have any loop-carried dependence
$\mathcal{P}$ the set of different possible pipeline configurations
$\mathcal{L}^{p}_{pip}$ the set of loops pipelined
$\mathcal{L}^{p}_{under\_pip_{l}}$ the set of loops under a loop pipelined
$\mathcal{L}^{p}_{above\_pip_{l}}$ the set of loops above the loop pipelined $l$
$\mathcal{L}_{s}$ the set of nested loops which iterate the statement $s$
$\mathcal{C}_{a_{d}}$ the set of loops which iterates the array $a$ at the dimension $d$
$d_{l}$ the maximum dependency distance of the loop $l$
Variable Description
$loop_{l}\_UF$ Unroll factor of the loop $l$
$loop_{l}\_tile$ TC of the innermost loop after strip-mining of the loop $l$
$loop_{l}\_pip$ Boolean to know if the loop $l$ is pipelined
$loop_{l}\_cache\_array\_a$ Boolean to know the the array $a$ is transferred on-chip before the loop $l$
Constant Description
$TC_{l}$ Trip Count of the loop $l$
$II_{l}$ II of the loop $l$
$IL_{par}$ Iteration Latency of the operations whitout dependencies of the statement $s$
$IL_{red}$ Iteration Latency of the operations whit dependencies of the statement $s$
$DSP_{s_{op}}$ Number of DSP used for the statement $s$ for the operation $op$
$DSP_{available}$ Number of DSP available for the FPGA used
$MAX\_PARTITIONING$ Maximum array partitioning defined by the user or the DSE (cf. Section 6)
Table 4, summarize the sets, variable and constants we use.
5.2. Modeling Compiler Transformations
Pipeline Rules
Vitis HLS unrolls all loops under the pipelined loop. This implies that all loop $l$ under the pipelined loop must have $loop_{l}\_UF==TC_{l}$ (Eq. 15). Considering this constraint, it is important to note that for each statement, only one of the loops that iterate the statement can have a pragma pipeline (Eq. 5). If multiple pipelined loops were present, the loops beneath the first pipelined loop would be unrolled instead.
Memory Transfer Rules
Merlin automates the process of transferring data on-chip and applying array partitioning. The tool caches data on-chip and packs it in chunks of up to 512 bits, enabling efficient transfer speed. When the data is already present on-chip it can be reused provided that resource constraints are satisfied. The compiler caches on-chip the data only above the loop pipelined (Eq. 14).
Dependencies
Loop-carried dependencies are managed using constraints (Eq 8). If a loop has a dependency distance of $n$, this means that if we unroll with an unroll factor $uf>n$ this is equivalent to unrolling the loop with a factor $uf=n$ because the statements corresponding to the iteration $\{uf+1,\dots,n\}$ will be executed only after the first n statements are executed due to the dependency. Loop-independent data dependencies are managed at the objective function, as elaborated in Section 5.4.
Supplementary Rules
In addition, we add the constraints for the maximum unrolling (Eq. 10), the divisibility of the problem size of the unroll factors (Eq. 6), the tile size (Eq. 7) and the maximum array partitioning (Eq. 13). For this last, this corresponds to adding an upper bound to the product of the UF of all the loops which iterate the same array on different dimensions.
During our DSE, we can force the solution to be fine-grained. In this case we add a constraint where the loop above the loop pipelined has a UF of 1, i.e., for all loop $l$ above the pipelined loop $loop_{l}\_UF==1$ (Eq. 9). We also constrain the resources, modeling their sharing optimistically. We consider the number of DSPs (Eq. 11) and on-chip memory (Eq. 12) used. As the consumption of DSPs can be difficult to estimate due to resource sharing we utilize an optimistic estimate, which considers a perfect reuse/sharing: as soon as a computation unit is free, its resource can be reused.
5.3. Constraints
(1) $\forall l\in\mathcal{L},1\leq loop_{l}\_UF\leq TC_{l}$
(2) $\forall l\in\mathcal{L},1\leq loop_{l}\_tile\leq TC_{l}$
(3) $\forall l\in\mathcal{L},loop_{l}\_pip\in\{0,1\}$
(6) $\forall l\in\mathcal{L},loop_{l}\_UF\%TC_{l}==0$
(7) $\forall l\in\mathcal{L},loop_{l}\_tile\%TC_{l}==0$
(8) $\forall l\in\mathcal{L},\text{ if }dd_{l}>1,loop_{l}\_UF\leq d_{l}$
(9) $\forall l\in\mathcal{L},\forall l^{\prime}\in\mathcal{L}_{above\_l},loop_{l}\_pip*loop_{l^{\prime}}\_UF<=1$
(10) $\forall s\in\mathcal{S},\prod_{l\in{L}_{s}}loop_{l}\_UF\leq MAX\_PARTITIONING$
(12) $\begin{cases}&\sum_{a\in\mathcal{A}}\sum_{l\in\mathcal{L}}loop_{l}\_cache\_array_{a}\\ &\times footprint\_array_{a}\_loop_{l}\leq Mem\end{cases}$
(13) $\begin{cases}&\forall a\in\mathcal{A},\forall(d,d^{\prime})\in\mathbb{N}^{2}\text{ with }d\neq d^{\prime},\forall l\in C_{a_{d}},\forall l^{\prime}\in C_{a_{d^{\prime}}},\\ &loop_{l}\_UF\times loop_{l^{\prime}}\_UF\leq MAX\_PARTITIONING\end{cases}$
(14) $\begin{cases}&\forall p\in\mathcal{P},\forall lp\in\mathcal{L}^{p}_{pip},\forall lbp\in\mathcal{L}_{under\_pip_{lp}}\forall a\in\mathcal{A},\\ &loop_{lbp}\_cache\_array_{a}==0\end{cases}$
(15) $\begin{cases}&\forall p\in\mathcal{P},\forall lp\in\mathcal{L}^{p}_{pip},\forall lbp\in\mathcal{L}_{under\_pip_{lp}},\\ &loop_{lp}\times loop_{lbp}\_UF==loop_{lp}\times TC_{lbp}\\ \end{cases}$
5.4. Objective function
In every loop nest, there will invariably be a pipeline loop due to either user-inserted or compiler-inserted instructions (AMD/Xilinx Merlin and Vitis automatically insert the pragma pipeline if it is not done by the user or the previous compiler). Therefore, the objective function takes the following form: $TC_{ap}\times(IL+II\times(\frac{TC}{UF}-1))$, where $TC_{ap}$ includes the loops situated above the pipeline. Parallel loops above the pipeline can be coarse-grained parallelized. The iteration latency within the unrolled loop body is divided into either reduction operations or non-reduction operations, as reduction operations require logarithmic time for the reduction process. The variable IL encompasses the latencies of the statements found within the pipelined loop body. Independent statements can be executed in parallel and statements with dependencies are summed, as detailed in Section 4.1.
5.5. Example
1Loop0: for(i=0; i<2100; i++)
2 S0:y[i] = 0;
3Loop1: for(i=0; i<1900; i++) {
4 S1:t[i] = 0;
5 Loop2: for(j=0; j<2100; j++)
6 S2:t[i]+=A[i][j]*x[j];
7 Loop3: for(j=0; j<2100; j++)
8 S3:y[j]+=A[i][j]*t[i];
9}
Listing 10: AtAx code for Large problem size: $t=A*x;y=A*t$
We now employ the AtAx kernel as an illustration. $S0$ and $S3$ do not have inter-iteration dependencies within their respective loops, namely $Loop0$ and $Loop3$. Therefore, it is possible to pipeline $Loop0$ and $Loop3$ with an $II\geq 1$. In other words, for all iterations within these loops, there are no dependencies on previous iterations within the same loop. Loop1 and Loop2 are reduction loops, and the reduction operation is an addition which has an IL of $IL_{+}$ cycles. So the II ¿= $IL_{+}$. If loop1 is pipelined, there is a dependency between S1, S2 and S3 Between S1 and S2 the distance is 1, so we just add IL_S1 and IL_S2. Between S2 and S3 the distance is log(N) so the final equation will be IL_S1 + IL_S2 * log(N) + IL_S3 for the loop body cycle. If statements can be run at the same time (i.e., there is no dependency) it is a max instead of an addition. If we encounter code like $for(j=2;j<N;j++)y[j]=y[j-2]+3;$, a straightforward approach to handling this type of dependency is to impose a constraint such as $loop_{l}\_UF\leq 2$, which is represented by equation 8. In this case, due to dependencies an $UF>2$ is similar to $UF=2$.
@@ 6. Design Space Exploration
We now present our Design Space Exploration (DSE) approach. Our approach focuses on identifying designs with the most promising theoretical latency within the available design space. However, it may result in suboptimal designs if the selected pragmas are not applied during compilation. To address this potential issue and ensure high QoR, we conduct an additional exploration within a restricted subspace. Our DSE explores two additional parameters: the type of parallelism and the maximum array partitioning factor. Array partitioning is a technique commonly used in FPGA contexts to divide arrays or matrices into smaller sub-arrays, which can be stored in independent memory blocks known as Block RAMs (BRAMs). AMD/Xilinx HLS has a limit of 1,024 partitions per array. The array partitioning is calculated by taking the product of loops that iterate the same arrays on different dimensions (cf. Section 5). So constraining the maximal array partitioning also constrains the maximal UF. This NLP based DSE technique is presented in the Algorithm 1. The DSE starts without constraint on parallelism and array partitioning. Then we alternate constraints on parallelism while decreasing the maximum unrolling factor and array partitioning.
@@ 7.1. Setup
We use kernels from Polybench/C 4.2.1 (PolyBench, [n. d.]). The complexities and sizes of the problems are detailed in Table 8 located in the Appendix. In addition we add a kernel of Convolution Neural Network (CNN). A single-precision floating point is utilized as the default data type in computations to compare to AutoDSE. Computations operate on medium and large datasets from PolyBench/C (PolyBench, [n. d.]) in order to have kernels with large footprint and have a large enough space to explore. Selecting medium and large problem sizes is crucial to accurately reflect the complexities encountered in various fields such as scientific simulations, data analytics, and artificial intelligence. These sizes pose challenges that mirror the practical limitations of memory transfer, where efficiently managing data movement becomes paramount due to its potential to bottleneck performance. In such contexts, the careful orchestration of memory transfers is essential to ensure optimal resource utilization and prevent computational inefficiencies. Additionally, the scale of these problems often exceeds on-chip memory capacity, necessitating strategies for effectively handling data footprints that surpass available memory, further emphasizing the need for meticulous memory management and optimization techniques. The problem size and loop order of CNN are J,I=256, P,Q=5 H,W=224. A description of each benchmark can be found in Tables 8 and 5. The ludcmp, deriche and nusinnov kernels are not present as PolyOpt-HLS (Pouchet et al., 2013a) does not handle negative loop stride. Cholesky and correlation contains a sqrt() operation which we do not support currently. Finally, we removed FDTD-2D because it exposed a bug in Merlin, and this generated a program where data dependencies are not fully preserved.
A double-precision floating point is utilized as the default data type in computations to compare to HARP. We chose to use the problem size used by HARP in order to reuse their model. The problem size and kernel use by HARP can be found on the Table 9.
@@
We analyze the kernels and automatically generate each NLP problem with a version of PolyOpt-HLS (Pouchet et al., 2013b). We modified and extended for our work. Employing the AMPL description language to solve the NLP problems, we ran the commercial BARON solver (Tawarmalani and Sahinidis, 2005; Sahinidis, 2017) version 21.1.13. For our experiments, we utilize 2 Intel(R) Xeon(R) CPU E5-2680 v4 @ 2.40GHz and 252GB DDR4 memory.
@@ 7.2.1. AutoDSE
We compare our method with AutoDSE (Sohrabizadeh et al., 2021), described in Section 2, and we automatically generate the space of AutoDSE with the command ds_generator. We replace the UF and tile size by all the UF and tile size which divide the TC in order to have the same space. AutoDSE does not impose any constraints on parallelism or the maximum array partitioning. It employs an incremental exploration approach, enabling it to make compiler-specific pragma selections.
@@ 7.3. Comparison with AutoDSE
Figures 2 and 3 show the comparison with AutoDSE for Large and Medium problem size respectively. Table 5 shows the details of the comparison with AutoDSE. NL, ND, S, and Space S are respectively the number of loops, the number of polyhedral dependencies (WaR, WaW, RaW), problem size (L for Large and M for Medium) and space size. For each method we compute the throughput (GF/s) in GFLOPs per second, the total time of the DSE (T) in minutes, the number of designs explored (DE) and the number of designs timeout (DT). In addition, for AutoDSE we add the number of design that are early rejected/prune (ER) as AutoDSE prunes the design when AMD/Xilinx Merlin can not apply one of the pragmas, due to its analysis limitations.
@@
gemver 7 13 L 1.28E+13 9.94 9.94 290 21 7 0.18 1,756 221 206 10 6.06x 54.69x
gesummv 2 17 M 6.12E+02 1.97 1.97 220 14 3 1.97 836 80 8 47 3.80x 1.00x
gesummv 2 17 L 6.33E+03 1.82 2.64 236 18 1 0.56 692 94 29 60 2.93x 4.73x
gramschmidt 6 34 M 1.75E+07 1.58 1.58 364 7 3 0.44 934 109 92 8 2.57x 3.56x
gramschmidt 6 34 L 1.22E+08 2.34 2.34 420 6 4 0.95 819 265 11 239 1.95x 2.47x
@@
floyd-warshall 3 21 M 8.65E+04 0.17 0.61 246 17 3 0.10 1,605 60 22 29 6.52x 6.15x
floyd-warshall 3 21 L 3.29E+06 0.17 1.31 381 20 6 0.10 2,728 150 71 77 7.16x 13.09x
@@
Table 5. Comparison of DSE time and Throughput for NLP-DSE , NLP-DSE-FS, and AutoDSE across Polybench kernels for different problem sizes
(a) GF/s
(b) DSE Time (min)
@@
For AtAx Large (Lst. 10), AutoDSE explores 166 designs of which 106 are early rejected and 30 timeout. AutoDSE starts by partially unrolling Loops 2 and 3 and will then attempt to do a coarse-grained parallelization on Loop 1 with all divisors, which is impossible due to dependencies. Although AutoDSE manages to prune/early reject the designs because Merlin cannot apply the pragmas, it still requires several minutes of compilation by Merlin for each unroll factor. In parallel, AutoDSE tries to pipeline the outermost loops (and therefore unroll the innermost loops) which creates numerous timeouts. Although the first two designs timeout due to too high level of parallelism, NLP-DSE allows us to find a configuration with the innermost loops unrolled with a UF=700. This allows us to find a design with a 3.46x higher throughput in 11.34x less time.
@@ 7.4. Comparison with HARP
Figure 4 shows the comparison with HARP for Small and Medium problem size. Table 9 in Appendix, shows the details of the comparison.
The throughput is 1.45x higher on average (1.20x for geo-mean) for the kernel evaluated in similar DSE time. For 20/23 kernels the method identifies a design with a throughput similar to (+/- 10%), or better than, HARP.
We note a variation in the enhancement of performance compared to the evaluation in Section 7.3 vs. AutoDSE, stemming from various factors. Primarily, the breadth of the exploration space plays a significant role. HARP has the capacity to traverse an average of 150,000 designs, enabling it to nearly exhaustively explore the entire space. Additionally, HARP is trained and/or fine-tuned with precise knowledge of the kernel and problem size, granting it deep insight into scenarios where pragmas are not applied and have enough training on these specific kernel to estimate the latency. This confers an advantage over AutoDSE, which treats the compiler as a black box.
@@
In the case of the medium-sized mvt kernel, a notable slowdown occurs because Merlin is transferring the array A twice. This array comprises 159,900 double-precision floats. The kernel is already constrained by memory bandwidth when transferring A once. The duration for a single transfer of A is 20,000 cycles, utilizing a bitwidth of 512 (maximum), amounting to a total of 40,000 cycles. Consequently, the kernel’s overall latency is 40,726 cycles. However, HARP successfully identified a configuration that enables Merlin to transfer A only once.

### v4

@@ 6.2. Summary
By composing all the theorems, this allows us to end up with the final latency lower bound of the program which is presented in theorems 6.20 for the computation and 6.21 for the computation and communication.
@@
Given available resource $DSP_{avail}$ and a program which contains a loop body $L$ with a computation latency $Lat^{L}_{computation}$ and a communication latency $Lat^{L}_{communication}$.
@@
$Lat_{L}=Lat^{L}_{computation}+Lat^{L}_{communication}$
@@ 7. Design Space Exploration
We now present our Design Space Exploration (DSE) approach. Our approach focuses on identifying designs with the most promising theoretical latency within the available design space. However, it may result in suboptimal designs if the selected pragmas are not applied during compilation. To address this potential issue and ensure high QoR, we conduct an additional exploration within a restricted subspace. Our DSE explores two additional parameters: the type of parallelism and the maximum array partitioning factor. Array partitioning is a technique commonly used in FPGA contexts to divide arrays or matrices into smaller sub-arrays, which can be stored in independent memory blocks known as Block RAMs (BRAMs). AMD/Xilinx HLS has a limit of 1,024 partitions per array. The array partitioning is calculated by taking the product of loops that iterate the same arrays on different dimensions (cf. Section 5). So constraining the maximal array partitioning also constrains the maximal UF. This NLP based DSE technique is presented in Algorithm 1. The DSE starts without constraint on parallelism and array partitioning. Then we alternate constraints on parallelism while decreasing the maximum unrolling factor and array partitioning.
@@ 8.1. Setup
We use kernels from Polybench/C 4.2.1 (Pouchet and Yuki, 2023). The complexities and sizes of the problems are detailed in Table 11 located in the Appendix. In addition, we add a layer of a Convolutional Neural Network (CNN) to demonstrate that our method works effectively on different types of kernels. A single-precision floating point is utilized as the default data type in computations to compare to AutoDSE (Sohrabizadeh et al., 2022b) and ScaleHLS (Ye et al., 2022). Computations operate on medium and large datasets from PolyBench/C (Pouchet and Yuki, 2023) in order to have kernels with large footprint and have a large enough space to explore. Selecting medium and large problem sizes is crucial to accurately reflect the complexities encountered in various fields such as scientific simulations, data analytics, and artificial intelligence. These sizes pose challenges that mirror the practical limitations of memory transfer, where efficiently managing data movement becomes paramount due to its potential to bottleneck performance. In such contexts, the careful orchestration of memory transfers is essential to ensure optimal resource utilization and prevent computational inefficiencies. Additionally, the scale of these problems often exceeds on-chip memory capacity, necessitating strategies for effectively handling data footprints that surpass available memory, further emphasizing the need for meticulous memory management and optimization techniques. The problem size and loop order of CNN are J,I=256, P,Q=5 H,W=224. A description of each benchmark can be found in Tables 11 and 6. The ludcmp, deriche and nusinnov kernels are not present as PolyOpt-HLS (Pouchet et al., 2013) does not handle negative loop stride. Cholesky and correlation contains a sqrt() operation which we do not support currently. Finally, we removed FDTD-2D because it exposed a bug in Merlin, and this generated a program where data dependencies are not fully preserved.
A double-precision floating point is utilized as the default data type in computations to compare to HARP. We chose to use the problem size used by HARP (small and medium) in order to reuse their model. The problem size and kernel use by HARP can be found on the Table 10.
@@
We analyze the kernels and automatically generate each NLP problem with a version of PolyOpt-HLS (Pouchet et al., 2013). We modified and extended for our work. Employing the AMPL description language to solve the NLP problems, we ran the commercial BARON solver (Tawarmalani and Sahinidis, 2005; Sahinidis, 2017) version 21.1.13. For our experiments, we utilize 2 Intel(R) Xeon(R) CPU E5-2680 v4 @ 2.40GHz and 252GB DDR4 memory.
@@ 8.2.1. AutoDSE
We compare our method with AutoDSE (Sohrabizadeh et al., 2022b), described in Section 2, and we automatically generate the space of AutoDSE with the command ds_generator. We replace the UF and tile size by all the UF and tile size which divide the TC in order to have the same space. AutoDSE does not impose any constraints on parallelism or the maximum array partitioning. It employs an incremental exploration approach, enabling it to make compiler-specific pragma selections.
@@ 8.3. Comparison with AutoDSE
Figures 1 and 2 show the comparison with AutoDSE for Large and Medium problem size respectively.
@@
(a) GF/s
(b) DSE Time (min)
Table 6 shows the details of the comparison with AutoDSE. NL, ND, S, and Space S are respectively the number of loops, the number of polyhedral dependencies (WaR, WaW, RaW), problem size (L for Large and M for Medium) and space size. For each method we compute the throughput (GF/s) in GFLOPs per second, the total time of the DSE (T) in minutes, the number of designs explored (DE) and the number of designs timeout (DT). In addition, for AutoDSE we add the number of design that are early rejected/prune (ER) as AutoDSE prunes the design when AMD/Xilinx Merlin can not apply one of the pragmas, due to its analysis limitations.
@@
gemver 7 13 L 1.28E+13 9.94 9.94 290 21 7 0.18 1,756 221 206 10 6.06x 54.7x
gesum. 2 17 M 6.12E+02 1.97 1.97 220 14 3 1.97 836 80 8 47 3.80x 1.00x
gesum. 2 17 L 6.33E+03 1.82 2.64 236 18 1 0.56 692 94 29 60 2.93x 4.73x
gram. 6 34 M 1.75E+07 1.58 1.58 364 7 3 0.44 934 109 92 8 2.57x 3.56x
gram. 6 34 L 1.22E+08 2.34 2.34 420 6 4 0.95 819 265 11 239 1.95x 2.47x
@@
floyd-w 3 21 M 8.65E+04 0.17 0.61 246 17 3 0.10 1,605 60 22 29 6.52x 6.15x
floyd-w 3 21 L 3.29E+06 0.17 1.31 381 20 6 0.10 2,728 150 71 77 7.16x 13.1x
@@
Table 6. Comparison of DSE time and Throughput for NLP-DSE, NLP-DSE-FS, and AutoDSE across Polybench kernels at different problem sizes: NL, ND, S, and Space S denote the number of loops, dependencies, problem size (L for Large and M for Medium), and space size, respectively. GF/s indicates throughput, T represents DSE time (min), DE signifies the number of explored designs, and DT denotes timeout designs. Additionally, ET reflects the count of designs early rejected/pruned by AutoDSE
@@
For AtAx Large (Listing 3), AutoDSE explores 166 designs of which 106 are early rejected and 30 timeout. AutoDSE starts by partially unrolling Loops 2 and 3 and will then attempt to do a coarse-grained parallelization on Loop 1 with all divisors, which is impossible due to dependencies. Although AutoDSE manages to prune/early reject the designs because Merlin cannot apply the pragmas, it still requires several minutes of compilation by Merlin for each unroll factor. In parallel, AutoDSE tries to pipeline the outermost loops (and therefore unroll the innermost loops) which creates numerous timeouts. Although the first two designs timeout due to too high level of parallelism, NLP-DSE allows us to find a configuration with the innermost loops unrolled with a UF=700. This allows us to find a design with a 3.46x higher throughput in 11.34x less time.
@@ 8.4. Comparison with HARP
Figure 3 shows the comparison with HARP for Small and Medium problem size.
@@
Table 10 in Appendix, shows the details of the comparison.
The throughput is 1.45x higher on average (1.21x for geo-mean) for the kernel evaluated in similar DSE time. For 20/23 kernels the method identifies a design with a throughput similar to (+/- 10%), or better than, HARP.
We note a variation in the enhancement of performance compared to the evaluation in Section 8.3 vs. AutoDSE, stemming from various factors. Primarily, the breadth of the exploration space plays a significant role. HARP has the capacity to traverse an average of 75,000 designs, enabling it to nearly exhaustively explore the entire space. Additionally, HARP is trained and/or fine-tuned with precise knowledge of the kernel and problem size, granting it deep insight into scenarios where pragmas are not applied and have enough training on these specific kernel to estimate the latency. This confers an advantage over AutoDSE, which treats the compiler as a black box.
We observe two significant slowdown with the kernel mvt for medium problem size and gemm for small problem size.
The kernel mvt does two matrix vector multiplications where the same matrix $A$ is read for the two statements in reverse order. Even though our model assumes that the same array $A$ can be used for both statements with the same partitioning based on unrolling, Merlin transfers the array $A$ twice—once for each statement. However HARP find the design which allow the compiler to transfer only one times the array $A$. As we can see for the small size HARP is not able to find the configuration which allow Merlin to transfer $A$ one time and achieve to find a design with the same QoR as us. For our design the transfer of the two A take 97.2% of the total latency. So transferring only one time allow to achieve at least the same throughput as HARP.
For Gemm small, HARP leverages Merlin’s double-buffering to transfer the output efficiently. Without this optimization, which was not included in our design space, HARP achieves a throughput of 14.64 GFLOPS/s.
@@
8.5. Comparison with ScaleHLS
We evaluated the throughput (GFLOPS/s) of NLP-DSE against ScaleHLS (S-HLS) (Ye et al., 2022) using medium-sized Polybench kernels with single-precision floating-point operations. The results are summarized in Table 7.
To ensure a fair comparison, we considered two scenarios based on the memory model:
  • Comparison 1: Including Memory Transfers Since ScaleHLS uses the Vivado flow, which assumes that data reside on-chip, we manually added memory transfers to the ScaleHLS designs. We optimized the transfers using the maximum burst size allowed by the problem (e.g., 128 bits for an array of size 210) and synthesized the design using the Vitis flow. To further improve efficiency, we overlapped off-chip to on-chip data loads and on-chip to off-chip writes wherever possible.
  • Comparison 2: Excluding Memory Transfers For this scenario, we followed the methodology described in the ScaleHLS paper, where Vivado assumes all data are on-chip, eliminating memory transfer overhead. For a fair comparison, we extracted only the computation cycles from the Vitis flow report for our framework, isolating pure computation performance without memory overhead.
With Memory Without Memory
Kernel Th. S-HLS Th. NLP-DSE Perf. Imp. Th. S-HLS Th. NLP-DSE Perf. Imp.
Comparison 1 Comparison 1 Comparison 2 Comparison 2
2mm 23.61 117.48 4.98x 27.91 247.33 8.86x
3mm 19.52 138.73 7.11x 34.62 203.09 5.87x
Atax 1.43 1.96 1.37x 4.85 181.70 37.43x
Bicg 1.65 0.99 0.60x 9.03 287.07 31.79x
Gemm 38.48 105.18 2.73x 44.83 197.29 4.40x
Gesummv 1.66 1.97 1.19x 9.97 513.83 51.52x
Mvt 10.48 7.77 0.74x 31.87 298.51 9.37x
Symm 0.10 0.20 2.09x 0.07 0.20 3.09x
Syrk 0.44 1.32 2.99x 0.01 1.33 134.21x
Syr2k 0.02 1.74 92.54x 0.02 4.23 214.18x
Average 11.63x 50.07x
Geo-Mean 2.89x 20.26x
Table 7. Comparison of the throughput (GFLOPS/s) of NLP-DSE and ScaleHLS (S-HLS) for medium-sized Polybench kernels. In Case 1, we manually inserted memory transfers with the maximum burst size allowed by the problem size, then synthesized the designs using the Vitis flow. In Case 2, we followed the methodology from the ScaleHLS paper, running the ScaleHLS design with the Vivado flow, which assumes data is already on-chip. For NLP-DSE*, we extracted only the computation cycles from the Vitis report, excluding the cycles associated with off-chip memory transfers.
Although ScaleHLS can perform code transformations that our framework does not currently implement, the design space we explore for pragma insertion is far more comprehensive and fully explored within seconds or minutes using our NLP solver. ScaleHLS relies on a QoR estimator to evaluate each pragma configuration, which is fast (in the order of milliseconds), but it cannot exhaustively explore a large design space, similar to the limitations of tools like HARP. Unlike ScaleHLS, our DSE is not focused solely on identifying a theoretical solution quickly. Instead, our NLP solver efficiently explores the entire design space to identify a theoretical optimal solution in seconds or minutes. Furthermore, our approach incorporates a DSE process to avoid cases where pragmas are not applied as intended during synthesis. In contrast, ScaleHLS does not validate whether the selected pragmas are correctly implemented by the compiler, potentially leading to discrepancies between expected and actual performance.
The results demonstrate that NLP-DSE delivers superior QoR. When memory transfers are included in the ScaleHLS results (Comparison 1), our framework achieves an average speedup of 11.63x across the evaluated kernels, with a geometric mean of 2.89x. When isolating computation latency (Comparison 2), NLP-DSE achieves an average speedup of 50.07x, with a geometric mean of 20.26x.
We did observe slowdowns in memory-bound kernels like bicg and mvt. In these cases, our manually optimized memory transfers were more efficient than those handled by Merlin. For instance, in the bicg kernel, Merlin repeatedly transfers two large arrays, consuming 99.2% of the total latency. If these transfers were managed more efficiently, our framework could achieve a throughput of 1.96 GFLOPS/s. A similar inefficiency is observed in the mvt kernel.
Overall, NLP-DSE consistently outperforms ScaleHLS. Its memory management approach makes NLP-DSE more general and delivers better QoR by including on-chip memory transfers as part of the exploration space. Additionally, the theoretical design space is fully explored through the implementation of a cost model formulated as an NLP. A lightweight DSE is also employed to avoid cases where pragmas are not applied as expected by the compiler.